OpenStudy (cutiekawdfi):

help please <3 fan and medal 1. what is the simplest form of the quotient? ^4√810/^4√2 2. what is the simplest form of the quotient? √3x^12y^10/√5x^6y^3

Directrix (directrix):

^4√810/^4√2 I don't know how to read the ^ symbol in the context of this problem. Maybe this is the fourth root of 810 divided by the fourth root of 2. Is that it?

OpenStudy (mizuki):

345√

OpenStudy (mizuki):

that is number 1

OpenStudy (mizuki):

2. √3x^12y^10/√5x^6y^3 = √(3x^12y^10/5x^6y^3) = √[(3/5 × x^(12–6) y^(10 – 3)] = √[(3/5 × x^6 y^7] = x^3y^3√(3y/5) -----

Directrix (directrix):

>>>345√ I don't know what this answer is with a radical symbol with no radicand.